Reports 1-1 of 1 Clear search Modify search
On-call intervention (General)
Oncall-system - 2:10 Sunday 04 June 2023 (60437) Print this report
On-call intervention

The following report has been submitted to the On-call interface.

On-call events -> Injection System

Title: ISYS main chiller failure

Author(s): chiummo

Called at: 00:30, 04-06-2023, by: Alarm or monitoring system
Remote intervention: Started: 00:35, 04-06-2023; Ended: 00:45, 04-06-2023
On-site intervention: Started: 01:00, 04-06-2023; Ended: 02:15, 04-06-2023
Status: Action pending
Operator when issue resolved: None


DMS notified that the coolant temperature of the ISYS main chiller was way above its usual value (24C instead of 19C). By looking at the evolution of the temperatures of the appliances cooled by this chiller, it was clear that very soon the neovan amplifier safety would have tripped because of temperature of diode 4 reaching 35C. This happened actually in a few minutes.
On site inspection in the EERoom found the supposedly on-duty chiller ON but with blank display. No signs of liquid leakage. Pump was still working (flow of coolant was not decreased), so it is likely that the failure concerned the heat exchange part. A quick attempt was tried to switch off and then on again and this restored the chiller display, with apparent full functionality of the chiller. Of course this chiller was not trusted, so it was switched off and then the cooling circuit was swapped to the spare chiller.
The Neovan amplifier could be switched on again, PMC relocked and so the IMC.

With the spare chiller, we have anyway reduced performances of the cooling circuit. Coolant flow was measured in around 6l/min and is now 3.8l/min. Coolant temperature was around 19C and is now 21.3C.
Both these parameters affect the working temperatures of neovan head and neovan pumping diodes. We know that Neovan head increased temperature is a source of alignment drift at the input of the PMC.
To be monitored and possibly improved on monday.

* Note that any files attached to this report are available in the On-call interface.

Comments to this report:
chiummo - 2:12 Sunday 04 June 2023 (60438) Print this report

the missing plot

Images attached to this comment
melo, andreazzoli, ciardelli, spinicelli - 16:34 Friday 09 June 2023 (60500) Print this report

Today we have connected a new chiller to replace the faulty one. This chiller was returned from repair recently and was tested in Optics Lab for a few days. It has been put in motion (Fig. 1), and it is functioning as the main chiller at the moment. The behaviour of the chiller will be monitored for the next days.

Images attached to this comment
Search Help